Unpolarized light is sent through a linear polarizer. The intensity leaving the polarizer is 12.0 candelas. A second polarizer is placed after the first at an angle of 45.0°. What is the intensity after the second polarizer?
1
Expert's answer
2020-06-17T09:35:59-0400

From. Malus's Law for Polarization ,

"I=I_o\cos^2(\theta)"

Now, we have given,"I_o=12\:candelas,\theta=45^{\circ}"

Thus,

"I=12\cos^2(45^{\circ})=6\:candelas"